# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ 6 Discussion
6.1 Comparison and Connection with Related Complexity Measures
@@
    Relationship with Nielsen’s Geometric Complexity: Nielsen complexity [38], from the perspective of differential geometry, defines complexity as the length of the geodesic connecting an initial state to a target state on the manifold of the unitary group, where penalty factors are used to suppress evolution in "difficult" directions. RCC forms a profound complementary relationship with it. Regardless of how the penalty factors are chosen, our main theorem, Theorem 3.1, can provide a rigorous lower-bound anchor point for any geometric path, derived from the fundamental principles of information theory. In other words, the length of any geometric path, after being converted into the number of atomic instruction steps, must be greater than or equal to the lower limit set by the RCC of the final state. This allows RCC to serve as a universal tool for the calibration and sanity-checking of geometric complexity.
@@
    Relationship with Krylov Complexity: Krylov complexity [39] measures complexity by analyzing the expansion of an operator in the Lanczos/Krylov basis generated by the Hamiltonian and the initial operator, and is closely related to quantum chaos and operator growth. The two focus on different physical aspects: Krylov complexity reflects the operator spreading and chaotic properties under continuous-time evolution, whereas RCC is rooted in the computational framework of discrete gate models and the system’s structural priors. An interesting research direction is to explore the quantitative relationship between RCC and Krylov complexity in the specific setting where the energy eigenbasis is taken as the reference set $R$. We conjecture that for certain chaotic systems, the two may exhibit a monotonic or order-preserving correspondence, which is worthy of systematic numerical and analytical verification in the future.
@@
    Relationship with Resource Theories such as "Mana": Resource theories represented by "mana" [40] aim to quantify non-stabilizer states, a key resource for achieving universal quantum computation. RCC and this type of specific resource theory form a complementary "cost-benefit" perspective: a high "mana" value is often associated with the ability to perform complex computational tasks, while our work shows that preparing such a high-"mana" state must itself incur a corresponding complexity cost in terms of RCC. RCC provides a unified, information-theory-based "manufacturing cost" pricing for all these specific quantum resources.
6.2 Limitations and Open Questions
The rigor of the RCC framework is built upon several key assumptions, which also define its applicability boundaries and the open questions that remain to be explored. It is worth emphasizing that, within a fixed subspace $H_{R}$, all pure states have the same leading term for RCC, $\log(d_{R})/\log(\Gamma_{R})$. As an information-theoretic lower bound, RCC aims to set a baseline for resource consumption, rather than to distinguish between quantum states that have the same spectral distribution but different geometric or entanglement structures—a common feature of all state functions based on von Neumann entropy. The most central theoretical assumption lies in the strict commutativity of the reference set $R$, which makes it naturally suited for systems defined by Abelian group symmetries or stabilizer subgroups. For systems defined by non-Abelian group symmetries, a possible path for generalization is as follows: technically, one can decompose the Hilbert space into a direct sum of irreducible representations (irreps) according to group representation theory, and use the normalized equally-mixed state of each irrep block (or its energy-truncated version) as the reference state; concurrently, define the CPTP maps that preserve this algebraic structure (such as the conditional expectation onto the center of the algebra) as free operations. (Note: The complete formal construction of this generalization scheme, including the definition of its reference state and the preservation of core properties, has been detailed in Appendix F.1.) Within this framework, the monotonicity of RCC is expected to be preserved based on the data processing inequality, but its proof as a lower bound on circuit complexity would need to be reconstructed. Another core open question is the near-tightness of RCC as a lower bound. In Eq. (3.1), we proved that RCC is a rigorous lower bound for $C_{\mathrm{opt}}$, but this does not answer whether $C_{\mathrm{opt}}$ can also be bounded from above by RCC. If such an upper bound exists, it would imply that RCC is not just a theoretical baseline, but a benchmark with strong predictive power for the true complexity. We hereby formalize this as a testable core conjecture:
@@ 6.2 Limitations and Open Questions
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (6.1)
@@
6.3 Resolving the Paradox of Pure State Complexity through an Observational Window
@@
The core definition of the RCC framework, $C_{R}(\rho)=(\log d_{R}-S(\rho))/\log\Gamma_{R}$, while ensuring structural fairness, also introduces a profound paradox: from this global perspective, all pure states ($S(\rho)=0$)—whether they are simple product states or graph states with highly non-local entanglement—possess the same maximal leading term of complexity, $\log(d_{R})/\log(\Gamma_{R})$. This conclusion of "pure state degeneracy" severely contradicts both our physical intuition and practical computational experience, as it seems to completely erase the vastly different internal correlation structures of different quantum states, which are precisely what determine the true computational cost of preparing them. The resolution to this paradox is rooted in a more profound physical principle: complexity is not an intrinsic, absolute property of a quantum state, but rather a relational property that emerges from the interaction between an observer and the system. The structural complexity of a state is "latent" within its internal global correlations [41];to quantify it, one must actively "reveal" it through a physical probe. This "physical probe" is precisely the "observational window" $E_{\Xi}$ introduced for thermodynamic analysis in Sec. 5.4. This mathematical tool formalizes a physically realistic observer with finite capabilities. Such an observer cannot perceive the entire global pure state of the system at once, but can only interact with the system through a specific perspective limited by their capabilities (e.g., being able to perform only local measurements). The core insight is that when such a constrained observer examines a pure state with a complex internal structure, the state will appear as a mixed state within their "window" due to the inability to resolve its delicate global correlations, thus giving rise to an "emergent" non-zero entropy $S(E_{\Xi}(\rho))>0$. Crucially, the magnitude of this emergent entropy is a direct "fingerprint" of how the internal correlation structure of the pure state responds to that specific mode of observation. A window that is highly "aligned" with the state’s structure will reveal its latent complexity to the greatest extent; whereas a simple product state will remain simple under any local window, and its emergent entropy will be minimal. Therefore, by shifting the theoretical framework from the global $C_{R}(\rho)$ to the observation-window-dependent $C_{R,\Xi}(\rho)$, we not only solve the technical problem of pure state degeneracy but, more importantly, we elevate RCC into a more complete, dynamic theory. This theory acknowledges the central role of observation in defining physical reality [42] and provides a conceptually profound and practically feasible pathway for quantifying the structural complexity of different pure states. To this end, we further rigorize the concept of the observational window $E_{\Xi}$ from Sec. 5.4, defining it as a conditional expectation acting on $H_{R}$: $E_{\Xi}$ is a unital, completely positive and trace-preserving (CPTP), and idempotent quantum channel, corresponding to some von Neumann subalgebra $A_{\Xi}\subseteq\mathfrak{B}(H_{R})$, and satisfying the condition of compatibility with the reference vacuum, $E_{\Xi}(\sigma_{R})=\sigma_{R}$. This compatibility is an extension of the principle of structural fairness, ensuring that the observation process does not distort the zero point of complexity. Using the monotonicity of relative entropy, $D(\rho||\sigma_{R})\geq D(E_{\Xi}(\rho)||E_{\Xi}(\sigma_{R}))$, and the fact that $E_{\Xi}(\sigma_{R})=\sigma_{R}$, one can obtain $S(E_{\Xi}(\rho))\geq S(\rho)$, and thus $C_{R,\Xi}(\rho)\leq C_{R}(\rho)$. More importantly, this algebraic structure endows complexity with a hierarchical monotonicity: if the observational capability is enhanced, i.e., $A_{\Xi_{1}}\subseteq A_{\Xi_{2}}$, then the idempotency of the conditional expectation ensures $E_{\Xi_{1}}\circ E_{\Xi_{2}}=E_{\Xi_{1}}$, which in turn implies $S(E_{\Xi_{1}}(\rho))=S(E_{\Xi_{1}}(E_{\Xi_{2}}(\rho)))\geq S(E_{\Xi_{2}}(\rho))$, thereby yielding a well-behaved complexity hierarchy $C_{R,\Xi_{1}}(\rho)\leq C_{R,\Xi_{2}}(\rho)\leq C_{R}(\rho)$. To connect this theoretical framework with physical practice, we provide an operational lower bound, which forms the basis for experimentally probing the structural complexity of pure states.
@@
Proposition 6.2 (Windowed Pinching Lower Bound). Let $\{P_{i}\}\subset A_{\Xi}$ be a set of orthogonal projection operators forming a complete set on $H_{R}$ ($\sum_{i}P_{i}=I_{H_{R}}$), with ranks $r_{i}=\mathrm{Tr}(P_{i})$. Let $p_{\Xi,i}=\mathrm{Tr}[P_{i}E_{\Xi}(\rho)]$ be the probability of observing outcome $i$ under the window $\Xi$. Then the windowed RCC (this paper uniformly uses the natural logarithm, with entropy in units of nats, consistent with the convention in Sec. 5.4 and Appendix I) satisfies the following inequality:
$C_{R,\Xi}(\rho)\geq\frac{\ln d_{R}-H(\{p_{\Xi,i}\})-\sum_{i}p_{\Xi,i}\ln r_{i}}{\ln\Gamma_{R}}$ (6.2)
@@
Proof Outline. Let $\sigma=E_{\Xi}(\rho)$. Apply a pinching map $\Delta_{P}(\sigma)=\sum_{i}P_{i}\sigma P_{i}$ to $\sigma$. By the entropy-increasing property of pinching, $S(\sigma)\leq S(\Delta_{P}(\sigma))$. At the same time, the entropy of the pinched state satisfies the block-entropy decomposition $S(\Delta_{P}(\sigma))=H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}S(\sigma_{i})\leq H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}\ln r_{i}$, where $\sigma_{i}=P_{i}\sigma P_{i}/p_{\Xi,i}$. Substituting this chain of inequalities back into the definition of $C_{R,\Xi}(\rho)$, $(\ln d_{R}-S(\sigma))/\ln\Gamma_{R}$, yields Eq. (6.2). (For technical details on the entropy increase due to pinching and the block-entropy decomposition used in this proposition, see Ref. [22]; for the rank-one dephasing entropy bound and experimental protocol, see D; for the axiomatic background of windowed monotonicity and notational/unit conventions, see Appendix I.) The physical insight of Eq. (6.2) is particularly crucial: it provides an operational "lever" to break the complexity degeneracy of pure states under a finite observational capability. For any pure state, the upper bound of the leading term of its ideal global complexity $C_{R}(\rho)$ is $\ln(d_{R})/\ln(\Gamma_{R})$. However, through a fixed local observational window $E_{\Xi}$, the value of $S(E_{\Xi}(\rho))$ for different pure states will differ significantly due to their different entanglement structures, thus creating a gap in their $C_{R,\Xi}(\rho)$ values. The significance of Eq. (6.3) is that it transforms $C_{R,\Xi}(\rho)$ from a theoretical construct into a physical quantity with direct operational meaning. While the definition of $C_{R,\Xi}(\rho)$ itself already conceptually solves the pure state degeneracy problem through $S(E_{\Xi}(\rho))$, Eq. (6.3) provides, from a more practical direction, an operational lower bound determined entirely by the classical measurement outcomes $\{p_{\Xi,i}\}$. This forms a key bridge connecting abstract theory with laboratory reality and gives rise to a powerful scheme for actively probing the latent complexity within pure states, which we term the "Experimental Squeeze of the Complexity Spectrum." Its core logic is that the choice of the measurement basis $\{P_{i}\}$ reflects the experimenter’s "informed probing" of the system’s intrinsic structural complexity. For a highly structured pure state, such as a graph state defined by a stabilizer group $S$, if we "intelligently" choose a measurement basis that is "aligned" with its stabilizer generators, then according to stabilizer theory, the measurement outcomes will be highly deterministic ($\pm 1$). This will cause the probability distribution $\{p_{\Xi,i}\}$ to become extremely sparse, with its Shannon entropy $H$ approaching zero. According to Eq. (6.3), this will cause its measurable complexity lower bound $C_{R,\Xi}(\rho)$ to be "squeezed" upwards, approaching its theoretical upper limit $\ln(d_{R})/\ln(\Gamma_{R})$. Conversely, a simple product state, lacking intrinsic non-local correlations, will inevitably appear more random when probed with the same complex measurement basis, resulting in a larger $H$ and a lower bound. This sharp contrast reveals a profound principle: the structural complexity of a state can only be most fully revealed when it is probed by an observer who "understands" its structural complexity. The significance of this "experimental squeeze" scheme is twofold. First, it constructs a feasible verification path for testing the core open question of this paper, Conjecture 6.1 (the near-tightness of RCC). By finding the optimal measurement bases that maximize the complexity lower bound for different classes of pure states (such as graph states, tensor network states, etc.), we can systematically establish "certificates of tightness" for their complexity. Second, it elevates the RCC framework from a static spectral measure to a complete, observation-dependent dynamical theoretical framework, in which the operational process of "finding the optimal measurement basis" is itself equivalent to the core task of "decoding the structural complexity of the quantum state."
6.4 A Dual-Constraint Framework: RCC and the Exact Quantum Speed Limit

### v3

@@ VI Discussion
VI.1 Comparison and Connection with Related Complexity Measures
@@
    Relationship with Nielsen’s Geometric Complexity: Nielsen complexity [NielsenGeometry06], from the perspective of differential geometry, defines complexity as the length of the geodesic connecting an initial state to a target state on the manifold of the unitary group, where penalty factors are used to suppress evolution in "difficult" directions. RCC forms a profound complementary relationship with it. Regardless of how the penalty factors are chosen, our main theorem, Theorem 3.1, can provide a rigorous lower-bound anchor point for any geometric path, derived from the fundamental principles of information theory. In other words, the length of any geometric path, after being converted into the number of atomic instruction steps, must be greater than or equal to the lower limit set by the RCC of the final state. This allows RCC to serve as a universal tool for the calibration and sanity-checking of geometric complexity.
@@
    Relationship with Krylov Complexity: Krylov complexity [parker2019universal] measures complexity by analyzing the expansion of an operator in the Lanczos/Krylov basis generated by the Hamiltonian and the initial operator, and is closely related to quantum chaos and operator growth. The two focus on different physical aspects: Krylov complexity reflects the operator spreading and chaotic properties under continuous-time evolution, whereas RCC is rooted in the computational framework of discrete gate models and the system’s structural priors. An interesting research direction is to explore the quantitative relationship between RCC and Krylov complexity in the specific setting where the energy eigenbasis is taken as the reference set $R$. We conjecture that for certain chaotic systems, the two may exhibit a monotonic or order-preserving correspondence, which is worthy of systematic numerical and analytical verification in the future.
@@
    Relationship with Resource Theories such as "Mana": Resource theories represented by "mana" [Howard2014] aim to quantify non-stabilizer states, a key resource for achieving universal quantum computation. RCC and this type of specific resource theory form a complementary "cost-benefit" perspective: a high "mana" value is often associated with the ability to perform complex computational tasks, while our work shows that preparing such a high-"mana" state must itself incur a corresponding complexity cost in terms of RCC. RCC provides a unified, information-theory-based "manufacturing cost" pricing for all these specific quantum resources.
VI.2 Limitations and Open Questions
The rigor of the RCC framework is built upon several key assumptions, which also define its applicability boundaries and the open questions that remain to be explored. It is worth emphasizing that, within a fixed subspace $H_{R}$, all pure states have the same leading term for RCC, $\log(d_{R})/\log(\Gamma_{R})$. As an information-theoretic lower bound, RCC aims to set a baseline for resource consumption, rather than to distinguish between quantum states that have the same spectral distribution but different geometric or entanglement structures—a common feature of all state functions based on von Neumann entropy. The most central theoretical assumption lies in the strict commutativity of the reference set $R$, which makes it naturally suited for systems defined by Abelian group symmetries or stabilizer subgroups. For systems defined by non-Abelian group symmetries, a possible path for generalization is as follows: technically, one can decompose the Hilbert space into a direct sum of irreducible representations (irreps) according to group representation theory, and use the normalized equally-mixed state of each irrep block (or its energy-truncated version) as the reference state; concurrently, define the CPTP maps that preserve this algebraic structure (such as the conditional expectation onto the center of the algebra) as free operations. (Note: The complete formal construction of this generalization scheme, including the definition of its reference state and the preservation of core properties, has been detailed in Appendix F.1.) Within this framework, the monotonicity of RCC is expected to be preserved based on the data processing inequality, but its proof as a lower bound on circuit complexity would need to be reconstructed. Another core open question is the near-tightness of RCC as a lower bound. In Eq. (6), we proved that RCC is a rigorous lower bound for $C_{\mathrm{opt}}$, but this does not answer whether $C_{\mathrm{opt}}$ can also be bounded from above by RCC. If such an upper bound exists, it would imply that RCC is not just a theoretical baseline, but a benchmark with strong predictive power for the true complexity. We hereby formalize this as a testable core conjecture:
@@ VI.2 Limitations and Open Questions
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (20)
@@
VI.3 Resolving the Paradox of Pure State Complexity through an Observational Window
@@
The core definition of the RCC framework, $C_{R}(\rho)=(\log d_{R}-S(\rho))/\log\Gamma_{R}$, while ensuring structural fairness, also introduces a profound paradox: from this global perspective, all pure states ($S(\rho)=0$)—whether they are simple product states or graph states with highly non-local entanglement—possess the same maximal leading term of complexity, $\log(d_{R})/\log(\Gamma_{R})$. This result contrasts with the intuition that complexity should depend on internal correlation structures, as all global pure states share the same leading-order complexity in this measure.The resolution to this paradox is rooted in a more profound physical principle: complexity is not an intrinsic, absolute property of a quantum state, but rather a relational property that emerges from the interaction between an observer and the system. The structural complexity of a state is "latent" within its internal global correlations [Wheeler1989];to quantify it, one must actively "reveal" it through a physical probe. This "physical probe" is precisely the "observational window" $E_{\Xi}$ introduced for thermodynamic analysis in Sec. V.4. This mathematical tool formalizes a physically realistic observer with finite capabilities. Such an observer cannot perceive the entire global pure state of the system at once, but can only interact with the system through a specific perspective limited by their capabilities (e.g., being able to perform only local measurements). The core insight is that when such a constrained observer examines a pure state with a complex internal structure, the state will appear as a mixed state within their "window" due to the inability to resolve its delicate global correlations, thus giving rise to an "emergent" non-zero entropy $S(E_{\Xi}(\rho))>0$. Crucially, the magnitude of this emergent entropy is a direct "fingerprint" of how the internal correlation structure of the pure state responds to that specific mode of observation. A window that is highly "aligned" with the state’s structure will reveal its latent complexity to the greatest extent; whereas a simple product state will remain simple under any local window, and its emergent entropy will be minimal. Therefore, by shifting the theoretical framework from the global $C_{R}(\rho)$ to the observation-window-dependent $C_{R,\Xi}(\rho)$, we not only solve the technical problem of pure state degeneracy but, more importantly, we extend the RCC framework to include observation dependence. This approach incorporates the role of the observer in defining physical reality [Bohr1928] and provides a method for quantifying the structural complexity of pure states. To this end, we further rigorize the concept of the observational window $E_{\Xi}$ from Sec. 5.4, defining it as a conditional expectation acting on $H_{R}$: $E_{\Xi}$ is a unital, completely positive and trace-preserving (CPTP), and idempotent quantum channel, corresponding to some von Neumann subalgebra $A_{\Xi}\subseteq\mathfrak{B}(H_{R})$, and satisfying the condition of compatibility with the reference vacuum, $E_{\Xi}(\sigma_{R})=\sigma_{R}$. This compatibility is an extension of the principle of structural fairness, ensuring that the observation process does not distort the zero point of complexity. Using the monotonicity of relative entropy, $D(\rho||\sigma_{R})\geq D(E_{\Xi}(\rho)||E_{\Xi}(\sigma_{R}))$, and the fact that $E_{\Xi}(\sigma_{R})=\sigma_{R}$, one can obtain $S(E_{\Xi}(\rho))\geq S(\rho)$, and thus $C_{R,\Xi}(\rho)\leq C_{R}(\rho)$. More importantly, this algebraic structure endows complexity with a hierarchical monotonicity: if the observational capability is enhanced, i.e., $A_{\Xi_{1}}\subseteq A_{\Xi_{2}}$, then the idempotency of the conditional expectation ensures $E_{\Xi_{1}}\circ E_{\Xi_{2}}=E_{\Xi_{1}}$, which in turn implies $S(E_{\Xi_{1}}(\rho))=S(E_{\Xi_{1}}(E_{\Xi_{2}}(\rho)))\geq S(E_{\Xi_{2}}(\rho))$, thereby yielding a well-behaved complexity hierarchy $C_{R,\Xi_{1}}(\rho)\leq C_{R,\Xi_{2}}(\rho)\leq C_{R}(\rho)$. To connect this theoretical framework with physical practice, we provide an operational lower bound, which forms the basis for experimentally probing the structural complexity of pure states.
@@
Proposition 6.2 (Windowed Pinching Lower Bound). Let $\{P_{i}\}\subset A_{\Xi}$ be a set of orthogonal projection operators forming a complete set on $H_{R}$ ($\sum_{i}P_{i}=I_{H_{R}}$), with ranks $r_{i}=\mathrm{Tr}(P_{i})$. Let $p_{\Xi,i}=\mathrm{Tr}[P_{i}E_{\Xi}(\rho)]$ be the probability of observing outcome $i$ under the window $\Xi$. Then the windowed RCC (this paper uniformly uses the natural logarithm, with entropy in units of nats, consistent with the convention in Sec. V.4 and Appendix I) satisfies the following inequality:
$C_{R,\Xi}(\rho)\geq\frac{\ln d_{R}-H(\{p_{\Xi,i}\})-\sum_{i}p_{\Xi,i}\ln r_{i}}{\ln\Gamma_{R}}$ (21)
@@
Proof Outline. Let $\sigma=E_{\Xi}(\rho)$. Apply a pinching map $\Delta_{P}(\sigma)=\sum_{i}P_{i}\sigma P_{i}$ to $\sigma$. By the entropy-increasing property of pinching, $S(\sigma)\leq S(\Delta_{P}(\sigma))$. At the same time, the entropy of the pinched state satisfies the block-entropy decomposition $S(\Delta_{P}(\sigma))=H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}S(\sigma_{i})\leq H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}\ln r_{i}$, where $\sigma_{i}=P_{i}\sigma P_{i}/p_{\Xi,i}$. Substituting this chain of inequalities back into the definition of $C_{R,\Xi}(\rho)$, $(\ln d_{R}-S(\sigma))/\ln\Gamma_{R}$, yields Eq. (21). (For technical details on the entropy increase due to pinching and the block-entropy decomposition used in this proposition, see Ref. [NielsenChuang2010]; for the rank-one dephasing entropy bound and experimental protocol, see D; for the axiomatic background of windowed monotonicity and notational/unit conventions, see Appendix I.) The physical insight of Eq. (21) is particularly crucial: it provides an operational "lever" to break the complexity degeneracy of pure states under a finite observational capability. For any pure state, the upper bound of the leading term of its ideal global complexity $C_{R}(\rho)$ is $\ln(d_{R})/\ln(\Gamma_{R})$. However, through a fixed local observational window $E_{\Xi}$, the value of $S(E_{\Xi}(\rho))$ for different pure states will differ significantly due to their different entanglement structures, thus creating a gap in their $C_{R,\Xi}(\rho)$ values. The significance of Eq. (22) is that it transforms $C_{R,\Xi}(\rho)$ from a theoretical construct into a physical quantity with direct operational meaning. While the definition of $C_{R,\Xi}(\rho)$ itself already conceptually solves the pure state degeneracy problem through $S(E_{\Xi}(\rho))$, Eq. (22) provides, from a more practical direction, an operational lower bound determined entirely by the classical measurement outcomes $\{p_{\Xi,i}\}$. This forms a key bridge connecting abstract theory with laboratory reality and gives rise to a powerful scheme for actively probing the latent complexity within pure states, which we term the "Experimental Squeeze of the Complexity Spectrum." Its core logic is that the choice of the measurement basis $\{P_{i}\}$ reflects the experimenter’s "informed probing" of the system’s intrinsic structural complexity. For a highly structured pure state, such as a graph state defined by a stabilizer group $S$, if we "intelligently" choose a measurement basis that is "aligned" with its stabilizer generators, then according to stabilizer theory, the measurement outcomes will be highly deterministic ($\pm 1$). This will cause the probability distribution $\{p_{\Xi,i}\}$ to become extremely sparse, with its Shannon entropy $H$ approaching zero. According to Eq. (22), this will cause its measurable complexity lower bound $C_{R,\Xi}(\rho)$ to be "squeezed" upwards, approaching its theoretical upper limit $\ln(d_{R})/\ln(\Gamma_{R})$. Conversely, a simple product state, lacking intrinsic non-local correlations, will inevitably appear more random when probed with the same complex measurement basis, resulting in a larger $H$ and a lower bound. This sharp contrast highlights a general principle: the structural complexity of a state can only be most fully revealed when it is probed by an observer who "understands" its structural complexity. The significance of this "experimental squeeze" scheme is twofold. First, it constructs a feasible verification path for testing the core open question of this paper, Conjecture 6.1 (the near-tightness of RCC). By finding the optimal measurement bases that maximize the complexity lower bound for different classes of pure states (such as graph states, tensor network states, etc.), we can systematically establish "certificates of tightness" for their complexity. Second, it elevates the RCC framework from a static spectral measure to a dynamical framework that incorporates observational limitations, in which the operational process of "finding the optimal measurement basis" is itself equivalent to the core task of "decoding the structural complexity of the quantum state."
VI.4 A Dual-Constraint Framework: RCC and the Exact Quantum Speed Limit